In mark-recapture studies a certain number of individuals in a population are captured, tagged, and then released. The population is then sampled again, and the fraction of tagged individuals in the sample (That is, the fraction of individuals in the sample that have been ‘recaptured’) is used to estimate the population size. Suppose 4 moose out of a population of 20 are captured, tagged, & released. If 4 individuals from this population are then subsequntly captured, what is the probability that half of them will be tagged? (ENTER ANSWER AS A FRACTION).
Answer #1

X~ Bin (n=4, P5 % = 0.20) formula P(x=x) = () C1-p35.7 Ex 23 x-.n () - )in! = 1*2*3*.*.*n : P(x+2) = (4) (0.23 (0.23 2121